Your teacher is giving you a test worth 100 points containing 40 questions. There are 2-point and 4-point questions on the test. How many of each type of question are on the test?
Answer by ankor@dixie-net.com(22740)

:
Your teacher is giving you a test worth 100 points containing 40 questions.
There are 2-point and 4-point questions on the test
:
Let x = number of 2-pt questions; then (40-x) = number of 4-pt questions
:
There are 100 points on the test
2x + 4(40-x) = 100
2x + 160 - 4x = 100
2x - 4x = 100 - 160
-2x = -60
x = -60/-2
x = + 30 2-pt question
:
I'll let you figure out how many 4-pt questions